# AI for Tabular Data

### v1

@@ Abstract
Natural language interfaces to tabular data must handle ambiguities inherent to queries. Instead of treating ambiguity as a deficiency, we reframe it as a feature of cooperative interaction, where the responsibility of query specification is shared among the user and the system. We develop a principled framework distinguishing cooperative queries, i.e., queries that yield a resolvable interpretation, from uncooperative queries that cannot be resolved. Applying the framework to evaluations for tabular question answering and analysis, we analyze the queries in 15 popular datasets, and observe an uncontrolled mixing of query types neither adequate for evaluating a system’s execution accuracy nor for evaluating interpretation capabilities. Our framework and analysis of queries shifts the perspective from fixing ambiguity to embracing cooperation in resolving queries. This reflection enables more informed design and evaluation for natural language interfaces for tabular data, for which we outline implications and directions for future research.
@@ 1 Rethinking Ambiguity in Analytical Natural Language Queries
We argue for a shift in perspective on ambiguity: Instead of a problem to be fixed, ambiguity in user-system cooperation is an expression of a user’s understanding of language, and a signal about user intent and implicit reliance on division of labor. Based on this framing, we develop (1) a framework to characterize analytical queries through the lens of cooperative interaction, defining them by what a user makes explicit versus what a system must infer. From the framework, we derive (2) evaluation criteria and apply them to 15 benchmarks for natural language interfaces to tabular data, revealing systematic misalignments with open-domain requirements. Finally, we distill (3) broader implications for better design and evaluation of tabular data analysis systems.
@@ C.2 Query Ambiguity
Prompt C.3: Procedural Specification Analysis You are an expert data analyst and annotator. Your purpose is to meticulously analyze a user query to determine if it is sufficiently specified to be answerable in an open-domain insight extraction setting. For this to be possible, the query must be self-contained and unambiguous, providing enough detail to execute an analytical procedure without requiring further user input. Core Principle: Imagine an analyst has already located the perfect, unambiguous dataset for the query (e.g., they have the table of "all player statistics for the 2024-2025 English Premier League season"). Your job is to decide if the query gives them clear enough instructions on what to do with that data. If they would have to come back and ask you, "how should I calculate ’top’?" or "what do you mean by ’relationship’?", then the query is procedurally underspecified. For this assessment, you must assume the data has already been perfectly identified. Do not worry about whether entities (like "players"), timeframes, or domains are ambiguous. Your only job is to assess the clarity of the analytical steps requested. Focus only on the "what to do and how to do it," not the "what data." Data Specification Dimensions: You will analyze how the query specifies the analytical operations the user wants to perform on the data. 1. Task Specification (Binary: Specified / Underspecified) (Clarifying questions are typically: "What does ... mean?",...) This dimension evaluates the overall analytical action requested. (i.e., the user’s goal) Specified: The query clearly states an executable operation (e.g., list, count, rank, calculate the average, find the correlation). Underspecified: The query’s intent is vague (e.g., "tell me about," "information on", "relationship between"). 2. Scope Specification (Binary: Specified / Underspecified) (Clarifying questions are typically: "How to...?", "By what metric...?",...) This dimension evaluates the clarity of the calculations and transformations applied to the data (i.e., the method to achieve the goal). If ambiguity arises from which data column or entity to use (e.g., for filtering the data), this is a Data Specification issue (Part B), not a Scope Specification issue. Specified: The query does not require any calculations or analytical methods, or these methods are sufficiently specified: The query is a simple lookup that does not require complex analytical parameters (e.g., "List the names of drivers...", "What role did she play?"). The query requires some calculation or application of an analytical method, and relevant parameters for that calculation are clear or can be assumed by a standard, common-sense default (e.g., using a standard unweighted average, using a logical (time) unit). The explicit criteria include: * Ranking: The specific metric used for ordering (e.g., top 5 by revenue). * Grouping & Aggregation: The categories for grouping and the function to apply (e.g., the average salary per department). * Metric Calculation: The formula for any derived values (e.g., GDP per capita). * Analytical Model: The method for assessing a relationship (e.g., the correlation between...). Underspecified: The query is missing an important analytical parameter. Use this category only if a missing parameter represents a significant analytical choice that could materially change the query’s intent or result (e.g., the metric for "best," the regression method to use). Do not use it for instances where a reasonable default can be assumed without clarification. Query: {query}

### v2

@@ Abstract
Natural language interfaces to tabular data must handle ambiguities inherent to queries. Instead of treating ambiguity as a deficiency, we reframe it as a feature of cooperative interaction where users are intentional about the degree to which they specify queries. We develop a principled framework based on a shared responsibility of query specification between user and system, distinguishing unambiguous and ambiguous cooperative queries, which systems can resolve through reasonable inference, from uncooperative queries that cannot be resolved. Applying the framework to evaluations for tabular question answering and analysis, we analyze the queries in 15 popular datasets, and observe an uncontrolled mixing of query types neither adequate for evaluating a system’s execution accuracy nor for evaluating interpretation capabilities. This conceptualization around cooperation in resolving queries informs how to design and evaluate natural language interfaces for tabular data analysis, for which we distill concrete directions for future research and broader implications.
@@ 1 Rethinking Ambiguity in Analytical Natural Language Queries
We argue for a shift in perspective: Instead of a problem to be fixed, ambiguity in user-system cooperation is an expression of a user’s understanding of language, and a signal about user intent and implicit reliance on division of labor. Based on this framing, we develop (1) a framework to characterize analytical queries through the lens of cooperative interaction, defining them by what a user makes explicit versus what a system must infer. From the framework, we derive (2) evaluation criteria and apply them to 15 benchmarks for natural language interfaces to tabular data, revealing systematic misalignments with open-domain requirements. Finally, we distill (3) broader implications for better design and evaluation of tabular data analysis systems.
@@ C.2 Query Ambiguity
Prompt C.3: Procedural Specification Analysis You are an expert data analyst and annotator. Your purpose is to meticulously analyze a user query to determine if it is sufficiently specified to be answerable in an open-domain insight extraction setting. For this to be possible, the query must be self-contained and unambiguous, providing enough detail to execute an analytical procedure without requiring further user input. Core Principle: Imagine an analyst has already located the perfect, unambiguous dataset for the query (e.g., they have the table of "all player statistics for the 2024-2025 English Premier League season"). Your job is to decide if the query gives them clear enough instructions on what to do with that data. If they would have to come back and ask you, "how should I calculate ’top’?" or "what do you mean by ’relationship’?", then the query is procedurally underspecified. For this assessment, you must assume the data has already been perfectly identified. Do not worry about whether entities (like "players"), timeframes, or domains are ambiguous. Your only job is to assess the clarity of the analytical steps requested. Focus only on the "what to do and how to do it," not the "what data." Procedural Specification Dimensions: You will analyze how the query specifies the analytical operations the user wants to perform on the data. 1. Intent Specification (Binary: Specified / Underspecified) (Clarifying questions are typically: "What does ... mean?",...) This dimension evaluates the overall analytical action requested. (i.e., the user’s goal) Specified: The query clearly states an executable operation (e.g., list, count, rank, calculate the average, find the correlation). Underspecified: The query’s intent is vague (e.g., "tell me about," "information on", "relationship between"). 2. Methodological Specification (Binary: Specified / Underspecified) (Clarifying questions are typically: "How to...?", "By what metric...?",...) This dimension evaluates the clarity of the calculations and transformations applied to the data (i.e., the method to achieve the goal). If ambiguity arises from which data column or entity to use (e.g., for filtering the data), this is a Data Specification issue (Part B), not a Scope Specification issue. Specified: The query does not require any calculations or analytical methods, or these methods are sufficiently specified: The query is a simple lookup that does not require complex analytical parameters (e.g., "List the names of drivers...", "What role did she play?"). The query requires some calculation or application of an analytical method, and relevant parameters for that calculation are clear or can be assumed by a standard, common-sense default (e.g., using a standard unweighted average, using a logical (time) unit). The explicit criteria include: * Ranking: The specific metric used for ordering (e.g., top 5 by revenue). * Grouping & Aggregation: The categories for grouping and the function to apply (e.g., the average salary per department). * Metric Calculation: The formula for any derived values (e.g., GDP per capita). * Analytical Model: The method for assessing a relationship (e.g., the correlation between...). Underspecified: The query is missing an important analytical parameter. Use this category only if a missing parameter represents a significant analytical choice that could materially change the query’s intent or result (e.g., the metric for "best," the regression method to use). Do not use it for instances where a reasonable default can be assumed without clarification. Query: {query}